OpenStudy (anonymous):

An orchestra consists of 25 members. The youngest member is 32 years old. She leaves and is replaced by a new member who is 30 years old. By how much does the replacement change the average age of the members of the orchestra?

OpenStudy (kropot72):

Let the original average age of the members = x Then the sum of the original ages = 25x The new sum of the ages = 25x - 32 + 30 = 25x -2 The new mean is therefore \[\frac{25x-2}{25}=x-0.08\] The replacement reduces the aveage age of the members by 0.08 year.

OpenStudy (kropot72):

average*